ElexusQuimby:

Find the following expression by multiplying. (a + b)^3

Avadpfwm:

x3 + 8y3 = 144

ElexusQuimby:

@avadpfwm wrote:
x3 + 8y3 = 144
What?

Avadpfwm:

@elexusquimby wrote:
@avadpfwm wrote:
x3 + 8y3 = 144
What?
144 is ur answer

ElexusQuimby:

Thanks. But i need the equation written out.

jhonyy9:

(a+b)^3 = (a+b)(a+b)(a+b) = ?

ElexusQuimby:

Thanks This is what i see

ElexusQuimby:

This is the answer I got. Somehow some way Idk I have no clue what I am doing by the way.